OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5^x+14=26

OpenStudy (anonymous):

5^x = 26-14 = 12 now take logs of both sides

OpenStudy (anonymous):

ln 5^x = ln 12 x ln5 = ln 12 the rest is easy

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

all you need do now is divide both sides by ln 5 giving x = ln12 / ln 5 just use ur calculator